# AIMC-Spec: A Benchmark Dataset for Automatic Intrapulse Modulation Classification under Variable Noise Conditions

### v1

@@ Abstract
A lack of standardized datasets has long hindered progress in automatic intrapulse modulation classification (AIMC)—a critical task in radar signal analysis for electronic support systems, particularly under noisy or degraded conditions. AIMC seeks to identify the modulation type embedded within a single radar pulse from its complex in-phase and quadrature (I/Q) representation, enabling automated interpretation of intrapulse structure. This paper introduces AIMC-Spec, a comprehensive synthetic dataset for spectrogram-based image classification, encompassing 33 modulation types across 13 signal-to-noise ratio (SNR) levels. To benchmark AIMC-Spec, five representative deep learning algorithms—ranging from lightweight CNNs and denoising architectures to transformer-based networks—were re-implemented and evaluated under a unified input format. The results reveal significant performance variation, with frequency-modulated (FM) signals classified more reliably than phase or hybrid types, particularly at low SNRs. A focused FM-only test further highlights how modulation type and network architecture influence classifier robustness. AIMC-Spec establishes a reproducible baseline and provides a foundation for future research and standardization in the AIMC domain.
@@ I Introduction
This paper introduces a new synthetic dataset designed for benchmarking AIMC using spectrogram-based image classification methods. It includes a diverse collection of modulation types, including both radar and communication waveforms, and covers a wide range of SNR levels to support consistent and structured evaluation. The dataset is fully synthetic, allowing complete control over pulse parameters and noise levels while ensuring consistent ground-truth labels. This design isolates algorithmic performance from hardware-specific or environmental variables. To demonstrate the dataset’s practical utility, five representative deep learning algorithms from recent AIMC literature were selected, re-implemented, and benchmarked under varied signal conditions, enabling comparative analysis across architectures, modulation types, and noise levels. A focus of this work is on how different deep learning architectures respond to a common spectrogram representation of radar signals, rather than on comparing transform or preprocessing strategies. This framing ensures that observed performance differences reflect architectural design rather than input representation bias.
@@ I-A Contributions
  • This work addresses a major gap in the AIMC field by introducing a publicly available synthetic dataset, AIMC-Spec, designed specifically for benchmarking image-based classification models. It includes 33 distinct intrapulse modulations and standardized signal definitions to support consistent comparison across future studies.
  • Benchmarking results are provided using five diverse AIMC approaches, covering varied modeling strategies and preprocessing pipelines. These evaluations span 13 SNR levels, with an additional focused assessment on FM-only classification to isolate model performance under constrained conditions.
@@
Figure 1: Spectrograms of a 4-frequency shift-keying (4FSK) signal across all SNR levels in AIMC-Spec, ranging from +10 dB (top left) to –20 dB (bottom right). SNR initially decreases in 5 dB steps from +10 to 0 dB and then alters to 2 dB steps from 0 to –20 dB. The progression illustrates the degradation of time–frequency structure as noise increases.
@@ III Data Set Properties
AIMC-Spec is divided into 13 SNR levels, each containing 33 intrapulse modulations with 1,000 signals per modulation. Signals were generated via a Python script that automatically selected parameters relevant to each modulation type. This section outlines the key dataset attributes, including signal configurations, noise conditions, modulation types, and file organization.
@@ III-A Pulse and Modulation Parameter Selection
Each signal consists of 100,000 samples, corresponding to a simulation time of 1 ms, sampled at 100 MHz. The Python-based generator assigned parameters such as carrier frequency, pulse width, and time of arrival randomly within predefined bounds. The carrier frequency was selected randomly within a 50 MHz range and applied to modulate the baseband signal.
@@
LFM-BPSK $cellcolorblack \cellcolorblack$
2FSK-BPSK $cellcolorblack \cellcolorblack$
@@ III-B Signal to Noise Ratio (SNR)
As the dataset is synthetically generated, it lacks natural noise. To approximate real-world conditions, Additive White Gaussian Noise (AWGN) was applied after signal generation. AWGN was chosen because it effectively models thermal noise, which is commonly encountered in communication systems due to its random and uniformly distributed characteristics across all frequencies. The dataset includes 13 SNR levels: +10, +5, and a range from 0 to –20 dB in steps of 2. Specifically: +10, +5, 0, -2, -4, -6, -8, -10, -12, -14, -16, -18, -20 dB.
Noise was added directly to the raw pulse before spectrogram transformation, using a fixed noise power reference of 1 dB. An SNR of 0 dB signifies parity between signal and noise power. Figure 1 presents spectrograms of a representative signal between +10 and –20 dB, illustrating the effect of noise on visual clarity. While AIMC-Spec currently models only AWGN, future releases will incorporate additional channel effects–such as multipath propagation, fading, and pulse jitter–to emulated real-world radar conditions more closely.
@@ III-C Intrapulse Modulations
AIMC-Spec includes 33 intrapulse modulation types grouped into frequency (13), phase (17), and hybrid (3) categories. These modulation types also serve as the class labels for classification tasks. FM signals are often visually distinct in spectrograms due to frequency sweeps. In contrast, phase modulations (PM) and hybrid modulations (HM) produce more subtle patterns—phase shifts introduce discontinuities or faint texture shifts that are less salient to image-based classifiers.
@@ III-D Data Set Structure
The dataset is stored in 13 folders, each storing 33 binary .h5 files—one per intrapulse modulation—using Python’s h5py module. Each file contains 1,000 pulses per class with a single pulse including 100,000 I/Q samples along with associated metadata. Each .h5 file is approximately 1.18 GB in size, with each SNR folder being 37 GB, making the whole data set 482 GB.
@@ IV Benchmark Algorithms
A curated selection of five algorithms was chosen to benchmark AIMC-Spec, reflecting both their influence in the field and their architectural diversity. The set spans convolutional, denoising, U-Net-like, pretrained, and transformer architectures, ensuring that different feature-extraction paradigms are represented. Each has demonstrated strong performance in prior studies under varying SNR conditions. Collectively, these models form a balanced cross-section of contemporary AIMC research and provide insight into how distinct neural architectures handle a common spectrogram representation. Figure 2 shows the preprocessing outputs for each model on a 4FSK signal at 0 dB SNR, with architectural summaries provided in Appendix VII (Figures 710).
@@ V-B Implementation
All spectrograms were generated using the open-source cupy Python library to accelerate FFT computation on GPU. A consistent 80:20 train–test split was applied to each model. Original loss functions and optimizers were preserved as described in their respective source papers.
Training was performed on an NVIDIA GeForce RTX 4090 GPU using a batch size of 32. Each model trained for up to 200 epochs, with early stopping (patience of 50 epochs) to prevent overfitting. During training, loss histories were recorded for each run. Final model evaluations included overall classification accuracy and confusion matrix generation.
@@ V-C Experiments
Two core experimental setups were designed to evaluate model performance under controlled and variable noise conditions. The first setup trained and tested each model on the same SNR level (denoted as SNR($x$):SNR($x$)), assessing performance when the noise environment remains consistent. The second setup trained on a combined dataset spanning all SNR levels (+10+10 dB to 20-20 dB) and tested on individual levels (SNR(ALL):SNR($x$)), measuring each model’s capacity to generalize under unseen noise intensities.
Each experimental setup was repeated in two variants. The first included all 33 intrapulse modulation types to assess full-spectrum performance. The second restricted the task to a subset of 13 frequency modulations, chosen for their well-defined spectral features and higher separability. This subset reflects the more tractable conditions often used in prior AIMC studies, where models are frequently benchmarked on modulation classes that produce visually distinct patterns in spectrograms. These two variants enable both a rigorous full-task benchmark and a focused evaluation under favorable modulation conditions.
@@ V-D Results
LDC-Unet achieved the strongest performance across all evaluated conditions, including at –20 dB in both the SNR(ALL):SNR($x$) and SNR($x$):SNR($x$) configurations. On the full modulation set, it reached 41.84% and 41.06% accuracy, respectively, while on the frequency-only (FM) subset, it attained 90.81% and 90.46%. These results are illustrated in Figures 3, 4, 5 and 6. Performance was consistently higher in the FM-only tests, reflecting the reduced complexity of this subset compared to the full modulation classification task. Notably, LDC-Unet was the only model to maintain the top position across all experimental conditions, regardless of SNR configuration or modulation subset.
In the SNR(ALL):SNR($x$) full-modulation experiment, LDC-Unet’s average margin over the second-best model (ViT) was 11.27%, with a specific difference of 12.36 percentage points at –20 dB. On the FM subset of the same setup, the gap widened to 15.77 points. For the SNR($x$):SNR($x$) configuration, LDC-Unet led CDAE-DCNN by 16.03 percentage points on the full task and by 26.62 on FM. These differences indicate that LDC-Unet maintained a consistent and measurable advantage regardless of training condition or task type.
ViT ranked second in the SNR(ALL):SNR($x$) setup, reaching 43.18% at 0 dB on the full set and 91.96% on FM, as shown in Figures 4 and 6. However, its performance dropped in the SNR($x$):SNR($x$) setup, with 26.85% on the full task and 51.58% on FM at 0 dB, suggesting a greater reliance on larger training datasets.
CDAE-DCNN achieved the second-best results in the SNR($x$):SNR($x$) experiments, with 40.38% on the full task and 90.35% on FM at 0 dB, outperforming both ViT and LPI-Net in this limited-data context, as shown in Figures 3 and 5.
Across all models, FM classification yielded higher accuracy compared to the full modulation task. For example, in the SNR($x$):SNR($x$) setup at –20 dB, LDC-Unet’s accuracy increased from 41.06% on the full modulation set to 90.46% on FM—a gain of over 49 percentage points. This large gap reinforces the relative ease of distinguishing frequency-only modulations and highlights the added challenge introduced by phase and hybrid types, particularly under high-noise conditions.
@@
LDC-Unet consistently reported the highest accuracy across SNR levels in nearly all experimental setups. CDAE-DCNN reported the second-highest accuracy in both the full and FM tasks under the SNR($x$):SNR($x$) setting, while ViT achieved second-best results in the more data-rich SNR(ALL):SNR($x$) configuration. STFT-CNN showed relatively strong accuracy considering its compact architecture, demonstrating that competitive accuracy does not strictly require large parameter counts—particularly when paired with well-structured time-frequency inputs.
A notable trend was the disparity between model performance on frequency modulations (FM) versus the full 33-class task. While some models, such as CDAE-DCNN and LPI-Net, saw relatively modest gains, others—particularly ViT—showed substantial improvement. This suggests that ViT’s architecture benefits more from clean, distinct modulation classes and especially from larger datasets. Given that the SNR(ALL) training setup yields a 13-fold increase in data volume (due to 13 SNR levels), ViT’s improvement in that setting aligns with known transformer characteristics: large data quantities are essential for optimal performance.
Figure 5: Accuracy (%) vs SNR (dB) for experiment SNR($x$):SNR($x$), using only frequency modulations.
Figure 6: Accuracy (%) vs SNR (dB) for experiment SNR(ALL):SNR($x$), using only frequency modulations.
@@
While overall accuracies provide a useful benchmark, deeper inspection of the accuracy–SNR trends in Tables IVVII reveals distinct behavioral patterns that align with each algorithm’s design.
LDC-Unet maintained the highest stability across all noise conditions, with less than a 30% accuracy drop between +10 dB and –20 dB. This resilience stems from its local dense skip connections and residual encoder blocks, which preserve multi-scale features and reconstruct fine-grained spectrogram detail even when noise dominates the input. The use of SSR-loss further enforces intra-class compactness and inter-class separation, allowing the model to maintain discriminative boundaries under low-SNR degradation.
CDAE-DCNN also performed strongly at low SNRs due to its denoising pre-training stage. However, compared with LDC-Unet, it exhibited a steeper decline below –10 dB, consistent with the CDAE’s tendency to over-smooth high-frequency or phase-rich modulations. This explains its slight disadvantage in the full-modulation task despite high FM-only accuracy.
ViT achieved high accuracy in the data-rich SNR(ALL):SNR($x$) configuration but degraded rapidly in SNR($x$):SNR($x$), reflecting its dependency on large and diverse training sets. The transformer’s global-attention mechanism generalizes well when abundant samples capture the full SNR spectrum, but without that diversity its performance becomes unstable—a known limitation of transformer architectures in small-sample radar tasks.
STFT-CNN and LPI-Net demonstrated consistent but modest performance, limited mainly by their input representations. STFT-CNN’s small receptive field and compact RGB images constrain the capture of long-term spectral patterns, while LPI-Net’s grayscale preprocessing reduces feature richness for overlapping or phase-coded signals. Nevertheless, both models highlight that well-structured spectrogram inputs can achieve competitive results even with lightweight CNNs.
Across all algorithms, frequency-only (FM) tasks consistently exceeded full-modulation performance by 40–50 percentage points at low SNRs. This disparity indicates that the principal challenge lies in recognizing phase and hybrid modulations, where spectral cues are less visually separable and inter-class boundaries overlap under noise. The different slopes of each model’s accuracy–SNR curve underscore architectural sensitivity to these factors: denoising or skip-connected models degrade smoothly, while attention-based and shallow CNNs show abrupt losses once noise dominates feature variance.
Collectively, these observations suggest that model behavior on AIMC-Spec is driven less by absolute accuracy and more by how effectively each architecture mitigates information loss under noise and class imbalance—insight critical for guiding future multi-frame and multi-modal designs.
@@
The preprocessing choices and input formats—previously detailed for each algorithm—appeared to influence classification outcomes. Algorithms that preserved more visual detail through color or interpolation tended to achieve higher accuracy on more complex modulation types. For example, LDC-Unet and CDAE-DCNN were better able to retain phase and hybrid modulation characteristics, likely due to their use of RGB inputs and larger spectrogram resolution.
By contrast, LPI-Net’s grayscale preprocessing likely reduced feature richness, particularly for modulations with overlapping frequency or phase behaviors. STFT-CNN, though using RGB inputs, was constrained by its very small image size, limiting its spatial feature representation. ViT, which received phase spectrogram inputs in a single-channel format, achieved lower accuracy in the SNR($x$):SNR($x$) setting—again suggesting that the quantity and richness of training data have a disproportionate impact on transformer-based models compared to CNNs. These observations suggest that aligning input design with architectural characteristics may benefit performance in noisy environments.

### v2

@@ Abstract
A lack of standardized datasets has long hindered progress in automatic intrapulse modulation classification (AIMC)—a critical task in radar signal analysis for electronic support systems, particularly under noisy or degraded conditions. AIMC seeks to identify the modulation type embedded within a single radar pulse from its complex in-phase and quadrature (I/Q) representation, enabling automated interpretation of intrapulse structure. This paper introduces AIMC-Spec, a comprehensive synthetic dataset for spectrogram-based image classification, encompassing 30 modulation types across 5 signal-to-noise ratio (SNR) levels. To benchmark AIMC-Spec, five representative deep learning algorithms—ranging from lightweight CNNs and denoising architectures to transformer-based networks—were re-implemented and evaluated under a unified input format. The results reveal significant performance variation, with frequency-modulated (FM) signals classified more reliably than phase-modulated (PM) types, particularly at low SNRs. A focused FM-only test further highlights how modulation type and network architecture influence classifier robustness. AIMC-Spec establishes a reproducible baseline and provides a foundation for future research and standardization in the AIMC domain.
@@ I Introduction
The AIMC-Spec dataset was previously introduced as a synthetic benchmark for reproducible evaluation of intrapulse modulation classification algorithms [5]. While the initial release provided structured coverage of a broad range of modulation types and SNR conditions, its practical use revealed several limitations. In particular, the original storage format resulted in a very large dataset footprint, exceeding 500 GB for the full release, which hindered accessibility and rapid experimentation. This work presents an updated version of AIMC-Spec that addresses these limitations through a redesigned data storage pipeline, reducing the dataset size to under 30 GB while preserving full signal fidelity. In addition, the modulation set has been refined and the dataset generation process streamlined to improve usability, consistency, and reproducibility for AIMC benchmarking.
This paper presents the updated AIMC-Spec dataset as a synthetic benchmark for spectrogram-based AIMC using image classification methods. It includes a diverse collection of modulation types, including both radar and communication waveforms, and covers a wide range of SNR levels to support consistent and structured evaluation. The dataset is fully synthetic, allowing complete control over pulse parameters and noise levels while ensuring consistent ground-truth labels. This design isolates algorithmic performance from hardware-specific or environmental variables. To demonstrate the dataset’s practical utility, five representative deep learning algorithms from recent AIMC literature were selected, re-implemented, and benchmarked under varied signal conditions, enabling comparative analysis across architectures, modulation types, and noise levels. A focus of this work is on how different deep learning architectures respond to a common spectrogram representation of radar signals, rather than on comparing transform or preprocessing strategies. This framing ensures that observed performance differences reflect architectural design rather than input representation bias.
@@ I-A Contributions
  • This work addresses a major gap in the AIMC field by introducing a publicly available synthetic dataset, AIMC-Spec, designed specifically for benchmarking image-based classification models. It includes 30 distinct intrapulse modulations and standardized signal definitions to support consistent comparison across future studies.
  • Benchmarking results are provided using five diverse AIMC approaches, covering varied modeling strategies and preprocessing pipelines. These evaluations span 5 SNR levels, with an additional focused assessment on FM-only classification to isolate model performance under constrained conditions.
@@
Figure 1: Spectrograms of an LFM-up signal across the SNR levels used in AIMC-Spec, ranging from +6 dB (left) to –6 dB (right) in 3 dB steps. The progression illustrates the gradual degradation of the characteristic linear frequency sweep as the noise level increases.
@@ III Data Set Properties
AIMC-Spec is divided into 5 SNR levels, each containing 30 intrapulse modulations with 1,000 testing signals and 100,000 training per modulation. Signals were generated via a Python script that automatically selected parameters relevant to each modulation type. This section outlines the key dataset attributes, including signal configurations, noise conditions, modulation types, and file organization.
@@ III-A Pulse and Modulation Parameter Selection
Each signal sample contains only the pulse itself, with no pure noise-only segments stored. The pulse width (PW) is uniformly distributed between 0.5 $\mu{s}$ and 18 $\mu{s}$, with a resolution of 1ns1\;ns. Signals are sampled at a frequency of 50 MHz. The carrier (RF) frequency is uniformly distributed within a 50 MHz bandwidth centered at 9500 MHz, with a valid step size of 0.001 MHz (1 kHz resolution). The dataset is constructed with a uniform distribution across all valid parameter combinations. To enforce diversity and prevent over-representation of specific parameter pairs, each modulation type allows at most three repeated instances of any particular PW–RF combination.
@@
[…]
@@ III-B Signal to Noise Ratio (SNR)
As the dataset is synthetically generated, it lacks natural noise. To approximate real-world conditions, Additive White Gaussian Noise (AWGN) is applied dynamically during data loading through a custom Python dataset class, rather than being permanently stored with the generated signals. AWGN was selected as it models thermal noise, which is commonly encountered in communication and radar systems due to its random nature and uniform spectral distribution. The dataset includes five SNR levels ranging from +6 to –6 dB in steps of 3 dB, specifically: +6, +3, 0, –3, and –6 dB.
Noise is added directly to the raw complex I/Q pulse prior to spectrogram transformation. Unlike the previous implementation, the SNR calculation is based solely on the average power of the pulse itself, excluding any zero-padding or non-signal regions. This ensures that the specified SNR accurately reflects the true signal energy content. An SNR of 0 dB therefore represents equality between the average pulse power and the noise power. Figure 1 presents representative spectrograms across varying SNR levels, illustrating the progressive degradation in visual clarity as noise increases. While AIMC-Spec currently models only AWGN, future releases will incorporate additional channel effects—such as multipath propagation, fading, and pulse jitter—to more closely emulate operational radar environments.
@@ III-C Intrapulse Modulations
AIMC-Spec includes 30 intrapulse modulation types grouped into frequency (15) and phase (15) categories. These modulation types also serve as the class labels for classification tasks. FM signals are often visually distinct in spectrograms due to frequency sweeps. In contrast, PM types produce more subtle patterns—phase shifts introduce discontinuities or faint texture shifts that are less salient to image-based classifiers.
@@ III-D Data Set Structure
The dataset is distributed as two directories: a training directory and a testing directory. Each directory contains 30 binary .h5 files (one per intrapulse modulation), stored using Python’s h5py library. Each .h5 file contains a collection of signals, where each entry comprises a single pulse represented by complex I/Q samples together with associated metadata. The total size of the training directory is 28 GB, while the testing directory is 288 MB.
The dataset was originally intended to support a randomly generated training split alongside a fixed, pre-defined test split. However, due to constraints associated with releasing the full dataset generation pipeline, this approach was not feasible. To address this, the released training data were generated at a substantially larger scale than required for a standard 80:20 split. A deterministic sub-sampling procedure is therefore provided to construct user-defined train/validation partitions at run time. Specifically, for a given random seed and target ratio (e.g., 80:20), the loader samples disjoint subsets from each modulation file without replacement, yielding a reproducible split while preserving per-modulation balance.
@@ IV Benchmark Algorithms
A curated selection of five algorithms was chosen to benchmark AIMC-Spec, reflecting both their influence in the field and their architectural diversity. The set spans convolutional, denoising, U-Net-like, pretrained, and transformer architectures, ensuring that different feature-extraction paradigms are represented. Each has demonstrated strong performance in prior studies under varying SNR conditions. Collectively, these models form a balanced cross-section of contemporary AIMC research and provide insight into how distinct neural architectures handle a common spectrogram representation. Figure 2 shows the preprocessing outputs for each model on a 4FSK signal at 0 dB SNR, with architectural summaries provided in Appendix A (Figures 58).
@@ V-B Implementation
All spectrograms were generated using the open-source cupy Python library to accelerate FFT computation on GPU. During early experimentation, on-the-fly spectrogram generation was identified as a significant computational bottleneck, substantially increasing training time. To mitigate this, spectrograms were precomputed offline and stored on disk prior to model training. Memory-mapped (memmap) files were then used to efficiently stream spectrogram data during training, enabling scalable access without loading the full dataset into RAM. This approach significantly improved I/O efficiency and overall training throughput.
A consistent 80:20 train–test split was applied to each model. Original loss functions and optimizers were preserved as described in their respective source papers.
Training was performed on an NVIDIA GeForce RTX 4090 GPU using a batch size of 32. Each model trained for up to 50 epochs, with early stopping (patience of 10 epochs) to prevent overfitting. During training, loss histories were recorded for each run. Final model evaluations included overall classification accuracy and confusion matrix generation.
@@ V-C Experiments
Experiments were designed to evaluate model robustness under varying noise conditions. Each model was trained on a combined dataset spanning five SNR levels (+6, +3, 0, –3, and –6 dB) and subsequently evaluated at each individual SNR level. This setup assesses the model’s ability to generalize across different noise intensities while maintaining performance under specific operating conditions.
Each experimental configuration was conducted in two variants. The first included all 30 intrapulse modulation types within AIMC-Spec, consisting of 15 frequency-modulated and 15 phase-modulated signals, to evaluate performance across the complete classification task. The second restricted the task to the subset of 15 frequency-modulated signals, selected for their well-defined spectral characteristics and higher separability in spectrogram representations.
Together, these two variants provide both a comprehensive benchmark across the full modulation space and a focused assessment under more controlled classification conditions.
@@ V-D Results
LDC-Unet achieved the strongest performance across all evaluated conditions for both the full modulation classification task and the FM subset, as shown in Figures 4 and 4. On the full modulation set, it obtained a mean accuracy of 66.91%, with performance increasing from 57.37% at –6 dB to 74.29% at +6 dB. On the FM subset, the model achieved a mean accuracy of 87.95%, ranging from 81.24% at –6 dB to 91.30% at +6 dB. Across all SNR levels and both experimental tasks, LDC-Unet consistently maintained the highest accuracy.
The next group of models consisted of ViT, LPI-Net, and CDAE-DCNN, which followed the same overall ranking in both experiments. ViT achieved the second-best results, reaching a mean accuracy of 61.19% on the full modulation task and 84.07% on the FM subset, followed closely by LPI-Net with 59.72% and 83.05%, respectively. CDAE-DCNN ranked fourth overall, achieving 53.39% mean accuracy on the full task and 78.95% on the FM subset. Although these models exhibited steady improvements as SNR increased, their performance remained consistently below that of LDC-Unet across the evaluated noise levels.
The STFT-CNN baseline produced the lowest performance across both experiments. It achieved a mean accuracy of 32.62% on the full modulation task, ranging from 24.06% at –6 dB to 38.68% at +6 dB, and 65.80% on the FM subset, increasing from 51.56% to 73.01% across the same SNR range. Despite this improvement with increasing SNR, STFT-CNN remained substantially below the other models. Across all algorithms, accuracy increased steadily with SNR, and the relative ranking of models remained consistent, with LDC-Unet outperforming ViT, LPI-Net, CDAE-DCNN, and STFT-CNN throughout the evaluated range.
@@
LDC-Unet consistently achieved the highest accuracy across all evaluated SNR levels in both experimental tasks. Among the remaining models, ViT reported the second-highest overall accuracy, followed closely by LPI-Net and CDAE-DCNN, while STFT-CNN produced the lowest performance across the evaluated conditions. Despite its comparatively compact architecture, STFT-CNN still demonstrated that meaningful classification performance can be achieved with lightweight CNN models when paired with structured time–frequency representations.
Across all algorithms, accuracy increased steadily as SNR improved, with the most pronounced gains occurring between –6 dB and 0 dB. The relative ranking of models remained consistent across the evaluated noise levels, with LDC-Unet outperforming ViT, LPI-Net, CDAE-DCNN, and STFT-CNN throughout the entire range. Despite these improvements at higher SNRs, the overall accuracy levels reflect the challenging nature of the AIMC-Spec dataset, which includes a large number of modulation types with overlapping spectral characteristics and realistic noise conditions. This complexity introduces significant inter-class ambiguity, particularly for phase-modulated signals, making robust classification under low-SNR conditions substantially more difficult than in many previously reported AIMC benchmarks.
@@
While overall accuracies provide a useful benchmark, deeper inspection of the accuracy–SNR trends in Figures 4 and 4 reveals distinct behavioral patterns that align with each algorithm’s design.
LDC-Unet maintained the highest stability across all noise conditions, with performance remaining relatively consistent across the evaluated SNR range from –6 dB to +6 dB. This resilience stems from its local dense skip connections and residual encoder blocks, which preserve multi-scale features and reconstruct fine-grained spectrogram detail even when noise dominates the input. The use of SSR-loss further enforces intra-class compactness and inter-class separation, allowing the model to maintain discriminative boundaries under low-SNR degradation.
CDAE-DCNN also performed strongly at lower SNRs due to its denoising pre-training stage. However, compared with LDC-Unet, it exhibited a steeper performance decline as noise increased. This behavior is consistent with the CDAE’s tendency to over-smooth high-frequency or phase-rich modulations, which can reduce discriminative detail in complex spectrogram structures.
ViT achieved competitive performance but remained slightly below the CNN-based architectures across both experimental tasks. Transformer models rely heavily on large and diverse training datasets to learn robust global attention patterns. While this mechanism can capture long-range dependencies in spectrogram images, it is also more sensitive to noise-induced feature variability when training data diversity is limited.
STFT-CNN and LPI-Net demonstrated consistent but more modest performance, primarily limited by their input representations. STFT-CNN’s compact spectrogram images constrain the capture of long-term spectral structures, while LPI-Net’s grayscale preprocessing reduces feature richness for overlapping or phase-coded signals. Nevertheless, both models highlight that well-structured spectrogram inputs can still achieve reasonable classification performance using comparatively lightweight CNN architectures.
Across all algorithms, the FM-only task consistently produced higher accuracy than the full modulation classification task. This difference indicates that the principal challenge lies in recognizing phase and hybrid modulations, where spectral cues are less visually separable and inter-class boundaries overlap under noise. The different slopes of each model’s accuracy–SNR curve further highlight architectural sensitivity to these factors: models with skip connections or denoising mechanisms degrade more gradually, whereas architectures with more constrained feature extraction exhibit sharper performance declines as noise increases.
Collectively, these observations suggest that model behavior on AIMC-Spec is driven less by absolute accuracy and more by how effectively each architecture preserves discriminative information under noise. This insight is particularly relevant for guiding future multi-frame and multi-modal model designs, where temporal and cross-representation cues may further improve robustness.
@@
The preprocessing choices and input formats—previously detailed for each algorithm—appeared to influence classification outcomes. Algorithms that preserved more visual detail through color channels or higher-resolution spectrogram representations tended to achieve higher accuracy on more complex modulation types. For example, LDC-Unet and CDAE-DCNN were better able to retain phase and hybrid modulation characteristics, likely due to their use of RGB inputs and larger spectrogram resolutions.
By contrast, LPI-Net’s grayscale preprocessing likely reduced feature richness, particularly for modulations with overlapping frequency or phase behaviors. STFT-CNN, although using RGB inputs, was constrained by its small image size, limiting the spatial representation of spectral features. ViT, which received phase spectrogram inputs in a single-channel format, achieved slightly lower overall accuracy compared with the leading CNN-based models, suggesting that the quantity and richness of input information may play an important role in transformer-based radar classification tasks.
These observations suggest that aligning input representation with architectural characteristics may improve robustness, particularly in noisy environments where subtle spectral cues become more difficult to distinguish.